# Lepage Dehak EPITA Research Laboratory (LRE)France

### v1

@@ 2.1.2 DINO
DINO [8] adopts a self-distillation training framework in which a student is trained to predict the outputs of a teacher. The teacher’s weights are updated using an EMA of the student’s weights. A larger set of augmented utterances of different frame lengths is considered, resulting in four short (local) and two long (global) frames. All inputs are processed by the student, but only global views are processed by the teacher. The student and teacher projectors output an embedding normalized with a temperature-softmax. Centering and sharpening are applied to the teacher embeddings to prevent one dimension from prevailing while discouraging collapse to the uniform distribution. $\mathcal{L}_{\text{DINO}}$ is defined as:
@@ 2.2.1 Framework
To capture the unaltered audio patterns, SSPS introduces a reference frame $\hat{\bm{x}_{i}}$ sampled from $u_{i}$ using a longer audio segment and no data-augmentation. Additionally, two memory queues are implemented in the framework: $\hat{\bm{Q}}$ with size $(|\hat{\bm{Q}}|,D_{\text{repr}})$ for storing reference representations $\{\hat{\bm{y}}_{i}\}_{i\in\mathcal{I}}$, and $\bm{Q}^{\prime}$ with size $(|\bm{Q}^{\prime}|,D_{\text{emb}})$ for storing positive embeddings $\{\bm{z}^{\prime}_{i}\}_{i\in\mathcal{I}}$. After a pre-defined number of epochs, pseudo-positive embeddings are sampled from $\bm{Q}^{\prime}$. For SimCLR, the positive embedding $\bm{z}_{i}^{\prime}$ is replaced by $\bm{q}^{\prime}_{pos(i)}$. For DINO, the teacher embeddings $\bm{z}_{i,t}^{\prime}$ are replaced by $\bm{q}^{\prime}_{pos(i),t}$.
@@ 2.2.2 Pseudo-positives sampling
K-means clustering is performed on reference representations in $\hat{\bm{Q}}$ to group the utterances into $K$ clusters. Clustering is performed at each epoch to refine the quality of assignments as SSL representations improve. $c_{i}$ denotes the cluster index of the $i$-th utterance, and $\bm{m}_{k}$ represents the centroid for the $k$-th cluster. The proposed method considers the following techniques to determine the cluster $\hat{c}_{i}$ from which to sample a pseudo-positive.
@@ 3.4 Evaluation protocol
The scoring of each test trial is the cosine similarity of $l_{2}$-normalized representation pairs, derived from full-length utterances. The performance is reported in terms of EER and minDCF with $P_{target}=0.01$, $C_{\text{miss}}=1$, and $C_{\text{fa}}=1$.
@@ 4.1 Effect of SSL positive sampling on SV performance
subsectionSelection of SSPS hyperparameters
@@ 4.3 Visualization of speaker representations
To illustrate the improvement in class compactness, Figure 2 presents the t-SNE of 10 speaker representations from the test set with SSL and SSPS positive sampling techniques. SSPS allows for reducing intra-class variance by matching different recording conditions to the same speaker identity during the training. Note that same-speaker representations already far apart in the latent space are not considered to belong to the same speaker class when employing SSPS.

### v2

# Lepage Dehak nocounter]EPITA Research Laboratory (LRE)France

@@ 2.1.2 DINO
DINO [8] adopts a self-distillation training framework in which a student is trained to predict the outputs of a teacher. The teacher’s weights are updated using an EMA of the student’s weights. A larger set of augmented utterances of different lengths is considered, resulting in four short (local) and two long (global) segments. All inputs are processed by the student, but only global views are processed by the teacher. The student and teacher projectors output an embedding normalized with a temperature-softmax. Centering and sharpening are applied to the teacher embeddings to prevent one dimension from prevailing while discouraging collapse to the uniform distribution. $\mathcal{L}_{\text{DINO}}$ is defined as:
@@ 2.2.1 Framework
To capture the unaltered audio patterns, SSPS introduces a reference segment $\hat{\bm{x}_{i}}$ sampled from $u_{i}$ using a longer audio segment and no data-augmentation. Additionally, two memory queues are implemented in the framework: $\hat{\bm{Q}}$ with size $(|\hat{\bm{Q}}|,D_{\text{repr}})$ for storing reference representations $\{\hat{\bm{y}}_{i}\}_{i\in\mathcal{I}}$, and $\bm{Q}^{\prime}$ with size $(|\bm{Q}^{\prime}|,D_{\text{emb}})$ for storing positive embeddings $\{\bm{z}^{\prime}_{i}\}_{i\in\mathcal{I}}$. After a pre-defined number of standard SSL training epochs, SSPS is enabled and pseudo-positive embeddings are sampled from $\bm{Q}^{\prime}$ such that $\bm{z}_{i}^{\prime}$ is replaced by $\bm{q}^{\prime}_{pos(i)}$ in the previously defined SSL objective functions.
@@ 2.2.2 Pseudo-positives sampling
At the beginning of each SSPS epoch, k-means clustering is performed on reference representations in $\hat{\bm{Q}}$ to group utterances into $K$ clusters, allowing the assignments to be progressively refined as SSL representations improve. $c_{i}$ denotes the cluster index of the $i$-th utterance, and $\bm{m}_{k}$ represents the centroid for the $k$-th cluster. The proposed method considers the following techniques to determine the cluster $\hat{c}_{i}$ from which to sample a pseudo-positive.
@@ 3.4 Evaluation protocol
The scoring of each test trial is the cosine similarity of $l_{2}$-normalized representation pairs, derived from the full-length utterances. The performance is reported in terms of EER and minDCF with $P_{target}=0.01$, $C_{\text{miss}}=1$, and $C_{\text{fa}}=1$.
@@ 4.1 Effect of SSL positive sampling on SV performance
4.2 Selection of SSPS hyperparameters
@@ 4.4 Visualization of speaker representations
To illustrate the improvement in class compactness, Figure 2 presents the t-SNE of 10 speaker representations from the test set with SSL and SSPS positive sampling techniques. SSPS allows for reducing intra-class variance by matching different recording conditions to the same speaker identity during the training. Note that same-speaker representations already far apart in the latent space are not considered to belong to the same speaker class when employing SSPS, as detailed in [21].